# Dense $\mathbb{X}$ Retrieval: What Retrieval Granularity Should We Use?

### v2

@@ 5.1 Passage Retrieval Performance
We report our evaluation results in Table 2. We observe that retrieval by propositions outperforms retrieval by sentence or passage on most tasks for both unsupervised and supervised retrievers.
@@
With all dense retrievers tested, proposition-level retrieval consistently outperforms sentence and passage-level retrieval on average across the five datasets. With the unsupervised retrievers, i.e. SimCSE and Contriever, we see an averaged Recall@5 improvement of $+12.0$ and $+9.3$ (35.0% and 22.5% relative improvement) respectively over five datasets.
@@
With the supervised retrievers, proposition-level retrieval still shows an advantage on average, yet the sizes of improvements are smaller. We hypothesize that this is due to these retrievers being trained on query-passage pairs. For instance, with DPR and ANCE, which have been trained on NQ, TQA, WebQ, and SQuAD, we observe that proposition and sentence level retrieval perform slightly worse compared to passage level on three out of the four datasets, with the exception of SQuAD. As shown in Table 2, all supervised retrievers demonstrate comparable performance across three levels of retrieval granularity in NQ, TQA, and WebQ.
@@
However, on datasets that the retriever model has not seen during training, we observe that retrieval by proposition demonstrates a clear advantage. For instance, most notably on SQuAD or EntityQuestions, we observe that proposition-based retrieval significantly outperforms the other two granularities. We see 17-25% Recall@5 relative improvement on EntityQuestions with relatively weak retrievers like DPR and ANCE. Furthermore, the Recall@5 of retrieval by proposition on SQuAD improved most on TAS-B and GTR, with 10-16% relative improvements.
@@
SimCSE Passage 8.1 16.3 22.6 33.7 7.7 14.9 9.8 17.8 10.9 17.5 11.8 20.0
Sentence 10.1 18.0 27.2 37.2 9.6 15.6 17.3 24.8 13.0 19.8 15.4 23.1
Proposition 12.7 20.2 28.4 37.7 11.2 17.2 18.0 25.1 18.3 25.0 17.7 25.0
Contriever Passage 11.1 22.4 25.7 41.4 6.8 14.9 15.6 27.7 10.9 21.5 14.0 25.6
Sentence 13.8 23.9 30.5 44.2 9.1 17.2 22.6 32.8 12.2 22.2 17.6 28.1
Proposition 16.5 26.1 37.7 48.7 13.3 19.9 25.6 34.4 16.1 27.3 21.8 31.3
@@
DPR Passage 24.8 36.1 40.3 51.0 14.0 22.2 12.4 21.7 18.6 25.9 22.0 31.4
Sentence 27.6 35.9 44.6 52.8 16.3 23.7 18.6 26.1 21.8 28.2 25.8 33.3
Proposition 28.3 34.3 45.7 51.9 19.0 23.8 19.8 26.3 26.3 31.9 27.8 33.6
ANCE Passage 27.1 38.3 43.1 53.1 15.2 23.0 15.3 26.0 23.4 31.1 24.8 34.3
Sentence 30.1 37.3 47.0 54.7 16.6 23.8 22.9 30.5 25.9 32.0 28.5 35.7
Proposition 29.8 37.0 47.4 53.5 19.3 24.1 22.9 30.1 29.1 33.7 29.7 35.7
TAS-B Passage 21.1 33.9 39.3 50.5 13.1 20.7 23.9 34.6 30.9 37.3 25.7 35.4
Sentence 24.6 33.9 43.6 52.3 14.4 21.4 33.8 40.5 31.4 36.1 29.6 36.8
Proposition 26.6 34.0 44.9 51.8 18.1 23.7 34.2 38.9 34.2 37.8 31.6 37.2
GTR Passage 23.4 34.5 38.7 49.3 13.1 20.1 23.9 33.8 31.3 36.7 26.1 34.9
Sentence 26.8 35.1 43.9 52.2 15.9 21.6 35.6 41.3 31.3 35.1 30.7 37.1
Proposition 29.5 34.4 45.9 52.6 18.7 23.8 37.0 40.4 34.1 37.1 33.0 37.7
Table 3: Open-domain QA performance (EM = Exact Match) under retrieve-then-read setting where the number of retrieved words to the reader QA model is limited at $l=100$ or $500$. We use UnifedQA V2 Khashabi et al. (2022) as the reader model. The first $l$ words from the concatenated top retrieved text unit are feed as input to the reader model. Underline denotes cases where the training split of the target dataset was included in the training data of the dense retriever. In most cases, we see better QA performance with smaller retrieval units.
5.2 Retrieval by Proposition $\Rightarrow$ Better Cross-Task Generalization
Our results indicate that the advantage of retrieval by proposition becomes most visible in cross-task generalization settings. We observe that on SQuAD and EntityQuestions, retrieval by proposition brings more performance gain over retrieval by passage.
To better understand where the improvements can be attributed, we conduct an additional analysis on EntityQuestions. As EntityQuestions features questions targeting the properties of longer-tail entities, we study how the retrieval performance under three different granularities is affected by the popularity of the target entity in question, i.e. whether the entity appears frequently in Wikipedia or not. We estimate the popularity of each entity with the following method. Given the surface form of an entity, we use BM25 to retrieve the top-1000 relevant passages from Wikipedia. We use the number of occurrences of the entity in its top-1000 passages as an estimate of its popularity. With the 20,000 test queries in EntityQuestion, around 25% of the target entities have a popularity value of less or equal to 3.
Figure 3 shows the passage retrieval performance vs. the popularity of the target entity in each question. Across all 6 dense retrievers, we observe that retrieving by proposition shows a much larger advantage over retrieving by passage with questions targeting less common entities. As the popularity of entities increases, the performance gap decreases. Our findings indicate that the performance gain from retrieval by proposition can mostly be attributed to queries for long-tailed information. This echoes our observation that retrieval by proposition improves the cross-task generalization performance of dense retrievers.
6 Results: Open-Domain QA
In this section, we study how the choice of retrieval granularity affects downstream open-domain QA tasks. We show that retrieval by proposition leads to strong downstream QA performance in the retrieve-then-read setting, where the number of retrieved tokens for input to the reader QA model is capped at $l=100$ or $500$ words.
@@ 6 Results: Open-Domain QA
6.1 QA Performance
Table 3 shows the evaluation results. Across different retrievers, we observe higher QA performance in terms of the EM@$l$ metric on average when using propositions as the retrieval unit. The unsupervised retrievers, SimCSE and Contriever, demonstrate improvements of +5.9 and +7.8 in the EM@100 score (50% and 55% relative improvement), respectively. The supervised retrievers, DPR, ANCE, TAS-B, and GTR, improve +5.8, +4.9, +5.9, and +6.9 EM@100 (26%, 19%, 22%, 26% relative improvement), respectively. Similar to our observations from passage retrieval evaluations, we find retrieval by proposition becomes more beneficial to downstream QA performance when the retriever has not been trained on the target dataset. In other cases, retrieval by proposition still holds an advantage, but with a smaller margin on average.
6.2 Propositions $\Rightarrow$ Higher Density of Question-Related Information
@@
We show the results in Figure 4 and Figure 7 with $l$ ranging from 0 to 500 across all five datasets. For a fixed retrieved word budget, proposition retrieval demonstrates a higher success rate compared to sentence and passage retrieval methods. The most significant improvement of proposition retrieval over passage retrieval occurs within the range of 100-200 words, which corresponds to roughly 10 propositions, 5 sentences, or 2 passages. As the word count further increases, the recall rates of the three granularity converge since all question-relevant information is included in the retrieved text.
6.3 Error Case Study
@@
To understand the source of errors from each type of retrieval granularity, we present and discuss four typical examples of mistakes in Table 4. With each example, we show the question and its corresponding top-1 retrieved text unit by the GTR retriever across the three granularities.
@@
We observe that with passage-level retrieval, the ambiguity of an entity or its references presents a challenge for dense retrievers, which echoes findings from Min et al. (2020). For instance, in example Q1, the question asks for “Super Bowl 50”, but the retrieved passage and sentence refers to “Super Bowl 5”. In Example Q2, passage retrieval fails to identify the part referring to the correct “atomic number”. Instead, the top-1 retrieved passage mentions “atomic number” in a different and irrelevant context to the question. Retrieval by sentences can also have a similar problem as retrieval by passages like Example Q1. Also, retrieval by sentences faces another challenge of lacking context. In Example Q3, sentence-based retrieval fails as the correct sentence in the retrieved passage uses “it” to refer to the pericardial sac.
@@
Retrieval by propositions tackles the aforementioned problems by ensuring each retrieval unit contains one piece of fact only and necessary context is incorporated in the propositions. However, proposition-based retrieval faces challenges with questions that involve multi-hop reasoning over long-range textual analysis. In Example Q4, the retrieved passage separately describes the actor’s name and the character they portray. There is not a single proposition that entails both the question and the answer.
@@
7 Related Work
Recent works on dense retrievers typically adopt a dual-encoder architecture Yih et al. (2011); Reimers and Gurevych (2019); Karpukhin et al. (2020); Ni et al. (2022). With dual-encoders, each query and document is encoded into a low-dimensional feature vector respectively, and their relevance is measured by a non-parametric similarity function between the embedding vectors Mussmann and Ermon (2016). Due to the limited expressivity from the similarity function, dual encoder models often generalize poorly to new tasks with scarce training data Thakur et al. (2021). To this end, previous studies use techniques such as data augmentation Wang et al. (2022); Yu et al. (2023a); Izacard et al. (2022); Gao and Callan (2022); Lin et al. (2023); Dai et al. (2023), continual pre-training Chang et al. (2020); Sachan et al. (2021); Oguz et al. (2022), task-aware training Xin et al. (2022); Cheng et al. (2023), hybrid sparse-dense retrieval Luan et al. (2021); Chen et al. (2022) or mixed strategy retrieval Ma et al. (2022, 2023) and so on to improve cross-task generalization performance of dense retrievers.
The motivation of our work echoes in part with multi-vector retrieval, e.g. ColBERT Khattab and Zaharia (2020), DensePhrase Lee et al. (2021a, b), ME-BERT Luan et al. (2021), MVR Zhang et al. (2022), where the retrieval model learns to encode a candidate retrieval unit into multiple vectors to increase model expressivity and improve retrieval granularity Seo et al. (2019); Humeau et al. (2019). Our work instead focuses on the setting where we do not update the dense retriever model or its parameters. We show that segmenting the retrieval corpus into finer-grained units of proposition can be a simple and orthogonal strategy for improving the generalization of dual encoders dense retrievers at inference time.
The idea of using propositions as a unit of text representation dates back to the Pyramid method in summarization evaluation Nenkova and Passonneau (2004), where model generated summary is evaluated by each proposition. Proposition extraction from text has been a long-standing task in NLP, with earlier formulations focusing on a structured representation of propositions, e.g. Open Information Extraction Etzioni et al. (2008), Semantic Role Labeling Gildea and Jurafsky (2000). More recent studies have found success in extracting propositions in natural language form via few-shot prompting with large language models Min et al. (2023); Kamoi et al. (2023), or finetuning smaller compact-sized models Chen et al. (2023b).
Retrieve-then-read, or more broadly – retrieval augmented generation, has recently merged as a popular paradigm for open-domain question answering Lewis et al. (2021); Jiang et al. (2023); Asai et al. (2023). While earlier works provide up to the top 100 retrieved passages for the downstream reader Izacard and Grave (2021); Kedia et al. (2022), the amount of allowed context is significantly reduced when using recent large language models (e.g. top 10) Touvron et al. (2023); Yu et al. (2023b), due to their limited context window length and inability to reason over long context Liu et al. (2023). Recent efforts have tried to improve the quality of the reader context by filtering or compressing the retrieved documents Wang et al. (2023); Xu et al. (2023). Our work offers a new perspective by leveraging a new retrieval unit, the proposition that not only reduces the context length but also offers greater information density, effectively addressing the issue.
8 Conclusion
We propose the use of propositions as retrieval units for indexing corpus to improve dense retrieval performance at inference time. Through our experiments on five open-domain QA datasets with six different dense retrievers, we discovered that retrieval by proposition outperforms passage or sentence in both passage retrieval accuracy and downstream QA performance with a fixed retrieved word budget. We introduce FactoidWiki, an indexed version of the English Wikipedia dump, where text from 6 million pages is segmented into 250 million propositions. We hope that FactoidWiki, along with our findings in the paper, will facilitate future research on information retrieval.
Limitations
The scope of our current study on the granularity of retrieval corpus has the following limitations. (1) Retrieval Corpus – Our study only focus on Wikipedia as the retrieval corpus, due to the fact that most open-domain QA datasets adopts Wikipedia as the retrieval corpus. (2) Types of dense retrievers evaluated – In the current version of the paper, we only evaluate on 6 types of popular dense retrievers, most of which follow bi- or dual-encoder architecture. In future versions, we will include and discuss results on a broader range of dense retrievers. (3) Language – Our current study is limited to English Wikipedia only. We leave the exploration on other languages to future work.
Appendix A Retrieval Corpus Processing
[…]
Appendix B Training the Propositionizer
To compare the proposition generation performance of different models, we set up a development set and an evaluation metric. The development set contains an additional 1,000 pairs collected by GPT-4 using the same approach as the training set. We evaluated the quality of the predicted propositions by the F1 score of two sets of propositions. Motivated by the F1 score of two sets of tokens in BertScore, we designed the F1 score for two sets of propositions. Let $P=\{p_{1},...,p_{n}\}$ denote the set of labeled propositions and $\hat{P}=\{\hat{p}_{1},...,\hat{p}m\}$ the set of predicted propositions. We use $\mathrm{sim}(p_{i},\hat{p}j)$ to represent the similarity between two propositions. Theoretically, any text similarity metric can be used. We chose BertScore with roberta-large configuration as our $\mathrm{sim}$ function since we wanted our metric to reflect the semantic difference between propositions. We define
[…]
Appendix C Offline Indexing
Appendix D Retrievers Models
[…]
Appendix E Additional Results
In Section 5.2, we demonstrated the advantage of retrieval by proposition over retrieval by sentence, particularly as the population of the entity decreases in EQ. We used the occurrence in the top-1000 paragraphs retrieved by BM25 as a proxy for popularity, rather than counting the number of hyperlinks to the entity used in Sciavolino et al., 2021. Therefore, the trend in the performance versus popularity plot shows some differences (Figure 6) between our results and those in Sciavolino et al., 2021. For example, some entities are ambiguous (e.g., 1992, a TV series). In such cases, the occurrence of the surface form of the entity is large. Simultaneously, questions related to ambiguous entities are challenging to answer, leading to lower recall.
(a) Where was [X] born? Refer to caption (b) Who was [X] created by?
@@ Appendix E Additional Results
Figure 6: Document retrieval recall vs. the popularity of the target entity in each question from the EntityQuestions dataset. We display the performance of two relations.
@@
Passage normal-$\Rightarrow$ Propositions

### v3

@@ 5.1 Passage Retrieval Performance
We report our evaluation results in Table 3. We observe that retrieval by propositions outperforms retrieval by sentences or passages on most tasks for both unsupervised and supervised retrievers.
@@
With all dense retrievers tested, proposition-level retrieval consistently outperforms sentence and passage-level retrieval on average across the five datasets. With the unsupervised retrievers, i.e. SimCSE and Contriever, we see an averaged Recall@5 improvement of $+12.0$ and $+9.3$ (35.0% and 22.5% relative improvement) on five datasets.
@@
With the supervised retrievers, proposition-level retrieval still shows an advantage on average, yet the sizes of improvements are smaller. We hypothesize that this is due to these retrievers being trained on query-passage pairs. For instance, with DPR, which have been trained on NQ, TQA, WebQ, and SQuAD, we observe that proposition and sentence level retrieval perform slightly worse compared to passage level on three out of the four datasets, with the exception of SQuAD. As shown in Table 3, all supervised retrievers demonstrate comparable performance across three levels of retrieval granularity in NQ, TQA, and WebQ.
@@
However, on datasets that the retriever model has not seen during training, we observe that retrieval by proposition demonstrates a clear advantage. For instance, most notably on SQuAD or EntityQuestions, we observe that proposition-based retrieval significantly outperforms the other two granularities. We see 25% Recall@5 relative improvement on EntityQuestions with relatively weak retrievers like DPR. Furthermore, the Recall@5 of retrieval by proposition on SQuAD improved most on GTR, with 16% relative improvements.
5.2 Retrieval on Finer-grained Index $\Rightarrow$ Better Cross-Task Generalization
Our results show the advantage of retrieval on proposition-level index in cross-task generalization settings. We observe that on SQuAD and Entity Questions, retrieval on the proposition-level index brings more performance gain over the passage-level index and sentence-level index.
To better understand where the improvements can be attributed, we conduct an additional analysis on Entity Questions. As Entity Questions features questions targeting the properties of longer-tail entities, we study how the retrieval performance under three different granularities is affected by the occurance of the target entity in question, i.e. whether the entity appears frequently in Wikipedia or not. We estimate the frequency of each entity with the following method. Given the surface form of an entity, we use BM25 to retrieve the top 1000 relevant passages from Wikipedia. We use the number of occurrences of the entity in its relevant passages as an estimate of its frequency. With the 20,000 test queries, around 25% of the target entities have an frequency value of less or equal to 3.
Figure 3 shows the passage retrieval performance vs. the frequency of the target entity in each question. Across all four dense retrievers, we observe that retrieving by proposition shows a much larger advantage over retrieving by passages with questions targeting less common entities. As the frequency of entities increases, the performance gap decreases. Our findings indicate that the performance gain from retrieval by proposition can mostly be attributed to queries for long-tailed information. This echoes our observation that retrieval on proposition-level index improves the cross-task generalization performance of dense retrievers.
5.3 Higher Passage Recall $\Rightarrow$ Higher Downstream QA Accuracy
To further understand whether the passage retrieval on a finer-grained index achieves higher downstream QA performance, we extract the answer from the retrieved passage by a QA reader, Fusion-in-decoder. The results are shown in Table 4.
Retrieval by proposition-level index achieves the highest average exact match (EM) on all four retriever models. Apart from limited exceptions, the proposition-level index achieves the highest EM for most retrieval tasks and on most datasets. We observe that the trend of downstream QA performance is highly consistent with passage retrieval recall, suggesting higher passage recall implies better downstream QA performance.
@@ 5.3 Higher Passage Recall $\Rightarrow$ Higher Downstream QA Accuracy
SimCSE Passage 20.5 22.9 49.7 52.9 24.5 24.6 13.7 16.6 20.7 25.5 25.8 28.5
Sentence 21.1 24.3 52.1 54.2 24.2 26.1 17.7 21.5 22.9 28.3 27.6 30.9
Proposition 22.0 26.0 51.0 53.9 23.5 27.0 18.6 22.7 25.9 33.6 28.2 32.6
Contriever Passage 24.5 28.7 54.7 57.9 25.7 26.9 17.7 24.2 25.6 32.5 29.6 34.1
Sentence 25.0 30.2 56.3 59.2 26.8 29.2 22.5 28.1 26.1 34.1 31.3 36.2
Proposition 25.8 30.3 56.8 60.0 26.8 29.9 24.8 29.7 27.1 36.5 32.3 37.3
@@
DPR Passage 30.6 33.7 56.5 60.3 25.0 26.8 14.2 18.9 26.4 31.6 30.6 34.3
Sentence 32.5 34.1 58.3 61.7 25.4 28.0 17.6 22.1 29.8 35.6 32.7 36.3
Proposition 31.5 33.8 57.6 60.6 27.1 28.2 18.2 22.6 32.9 39.7 33.5 37.0
GTR Passage 30.0 33.9 56.9 60.0 24.5 25.9 21.5 27.4 42.2 45.3 35.0 38.5
Sentence 30.9 34.0 58.9 61.9 24.5 27.0 29.8 31.7 42.9 45.9 37.4 40.1
Proposition 32.1 33.8 58.8 62.3 25.7 29.1 32.5 33.1 43.0 48.1 38.4 41.3
Table 5: Open-domain QA performance (EM = Exact Match) with LLaMA-2-7B model (Touvron et al., 2023). The context in the prompts is constructed by passage, sentence, or propositions limiting at $l=100$ or $500$ tokens. We prompt the LLaMA-2-7B model with four-shot demonstrations for each test case.
@@
6 How Does Granularity Influence Retrieval-Augmented LMs?
In this section, we study how the choice of different granularity used in the prompts affects the retrieval-augmented generation across open-domain QA tasks. To fairly compare different granularity with the same computation budget, we limit the number of retrieved tokens for input to the language model at $l=100$ or $500$ tokens. Our results suggest that retrieval by finer-grained units enables a higher density of question-related information in the prompts, leading to better performance.
6.1 Open-domain QA Performance
Table 5 shows the evaluation results with LLaMA-2-7B as the language model. Across different retrievers, we observe higher QA performance in terms of the EM@$l$ metric on average when using propositions as the retrieval unit.
Using propositions rather than passages in the prompts, the four dense retrievers—SimCSE, ConRetriever, DPR, and GTR—improve by +4.1, +3.2, +2.7, and +2.8 in the EM@500 score. The improvements for using sentences over passages for the four retrieval models are +2.4, +2.1, +2, and +1.6, respectively. It is interesting to note that in the LLaMA-2-7B model, the QA accuracy on TQA and WebQ is not sensitive to retrieval type. The highest improvements over the closed-book setting are only +4.9 and +3.2, achieved by GTR with propositions. Nevertheless, we observe that using sentences and propositions in the prompts results in higher performance than using passages for all retrieval models on these two datasets. The results suggest that using finer-grained units in the prompts is beneficial to retrieval-augmented generation.
6.2 Finer-grained Granularity $\Rightarrow$ Higher Density of Question-Related Information
@@
We show the results in Figure 4 and 7 with $l$ ranging from 0 to 500 across all five datasets. For a fixed word retrieval budget, proposition retrieval shows a higher success rate than sentence and passage retrieval methods. The largest improvement of proposition retrieval over passage retrieval occurs within the range of 100-200 words, which corresponds to roughly 10 propositions, 5 sentences, or 2 passages. As word count increases, the recall rate of the three granularities converges, encompassing all relevant information.
7 Related Work
Recent works on dense retrievers typically adopt a dual-encoder architecture Yih et al. (2011); Reimers and Gurevych (2019); Karpukhin et al. (2020); Ni et al. (2022). With dual-encoders, each query and document is encoded into a low-dimensional feature vector respectively, and their relevance is measured by a non-parametric similarity function between the embedding vectors Mussmann and Ermon (2016). Due to the limited expressivity from the similarity function, dual encoder models often generalize poorly to new tasks with scarce training data Thakur et al. (2021). Previous studies use techniques such as data augmentation Wang et al. (2022); Yu et al. (2023a); Izacard et al. (2022); Gao and Callan (2022); Lin et al. (2023); Dai et al. (2023), continual pre-training Chang et al. (2020); Sachan et al. (2021); Oguz et al. (2022), task-aware training Xin et al. (2022); Cheng et al. (2023), hybrid sparse-dense retrieval Luan et al. (2021); Chen et al. (2022), or mixed strategy retrieval Ma et al. (2022, 2023) and so on to improve cross-task generalization performance of dense retrievers.
The motivation of our work echoes in part with multi-vector retrieval, e.g. ColBERT Khattab and Zaharia (2020), DensePhrase Lee et al. (2021a, b), ME-BERT Luan et al. (2021), and MVR Zhang et al. (2022), where the retrieval model learns to encode a candidate retrieval unit into multiple vectors to increase model expressivity and improve retrieval granularity Seo et al. (2019); Humeau et al. (2019). Our work instead focuses on the setting where we do not update the dense retriever model or its parameters. We show that indexing the retrieval corpus by different granularity can be a simple and orthogonal strategy for improving the generalization of dense retrievers at inference time.
In line with generating retrieval units from the original corpus, Sarthi et al. (2024) propose using generative summaries as additional retrieval units alongside the original text, enhancing queries with document-level understanding. In contrast, our work generates propositions to improve queries related to long-tailed entities. These approaches are complementary, as they address different aspects of retrieval enhancement.
The use of propositions as a unit of text representation dates back to the Pyramid method in summarization evaluation Nenkova and Passonneau (2004), where a model-generated summary is evaluated by each proposition. Proposition extraction from text has been a long-standing task, with earlier formulations focusing on a structured representation of propositions Etzioni et al. (2008); Gildea and Jurafsky (2000). More recent studies have found success in extracting free-text propositions via few-shot prompting with LLMs Min et al. (2023); Kamoi et al. (2023), or fine-tuning compact-sized models Chen et al. (2023b).
Retrieve-then-read, or more broadly retrieval augmented generation, has recently emerged as a popular paradigm for open-domain question answering Lewis et al. (2021); Jiang et al. (2023); Asai et al. (2023). While earlier works provide up to the top 100 retrieved passages for the downstream reader Izacard and Grave (2021); Kedia et al. (2022), the amount of context allowed is significantly reduced when using recent large language models Touvron et al. (2023); Yu et al. (2023b), due to the limited context window length and inability to reason over long context Liu et al. (2023). Recent efforts try to improve the quality of the reader context by filtering or compressing the retrieved documents Wang et al. (2023); Xu et al. (2023). Our work offers a new perspective by changing the retrieval granularity, in order to achiev greater information density with a fixed context length.
8 Conclusion
This paper studies how the choice of granularity for indexing a corpus, as well as the granularity used in the prompts, influences retrieval and downstream QA performance. Our results show that retrieval by propositions outperforms passage-level and sentence-level retrieval on passage retrieval and downstream QA across five open-domain QA datasets. Our analysis shows that indexing a corpus with finer-grained units enhances the cross-task generalization of dense retrievers and increases the density of question-related information in the prompts. We hope that FactoidWiki and our findings will facilitate future research on information retrieval and retrieval-augmented generation.
Limitations
The scope of our current study on the granularity of retrieval corpus has the following limitations. (1) Retrieval Corpus – Our study only focuses on Wikipedia as the retrieval corpus, due to the fact that most open-domain QA datasets adopt Wikipedia as the retrieval corpus. (2) Types of dense retrievers evaluated – In the current version of the paper, we evaluate 6 types of popular dense retrievers, most of which follow the bi- or dual-encoder architecture. In future versions, we will include and discuss results on a broader range of dense retrievers. (3) Language – Our current study is limited to English Wikipedia only. We leave the exploration on other languages to future work.
Ethical Considerations
This article follows the ACL Code of Ethics. Our work is a foundational research on information retrieval. To the best of our knowledge, we do not find obvious risks related to malicious harmful effects, environmental impact, fairness considerations, or privacy considerations.
Acknowledgements
The authors sincerely appreciate anonymous reviewers for helpful discussions and comments. The authors would like to thank Xuanyu Ben Zhou, Ruixin Hong, Ning Dai, and Linfeng Shen for valuable feedback on the project. Xinran Zhao is supported by the ONR Award N000142312840.
Appendix A Retrieval Corpus Processing
[…]
Appendix B Training the Propositionizer
To compare the proposition generation performance of different models, we set up a development set and an evaluation metric. The development set contains an additional 1,000 pairs collected by GPT-4 using the same approach as the training set. We evaluated the quality of the predicted propositions by the F1 score of two sets of propositions. Motivated by the F1 score of two sets of tokens in BertScore, we designed the F1 score for two sets of propositions. Let $P=\{p_{1},...,p_{n}\}$ denote the set of labeled propositions and $\hat{P}=\{\hat{p}_{1},...,\hat{p}m\}$ the set of predicted propositions. We use $\mathrm{sim}(p_{i},\hat{p}j)$ to represent the similarity between two propositions. Theoretically, any text similarity metric can be used. We chose BertScore (Zhang et al., 2020) with roberta-large Liu et al. (2020) configuration as our $\mathrm{sim}$ function since we wanted our metric to reflect the semantic difference between propositions. We define
[…]
Appendix C Quality Analysis of Generated Propositions
We collected propositions generated from 50 randomly selected passages. There are 408 and 445 propositions generated by GPT-4 and Propositionizer, respectively. The propositions and passages were provided to an expert without knowing which model generated each proposition. The expert annotated three scores from different perspectives for each proposition: (1) whether the proposition is fully supported by the passage, (2) whether the proposition is minimal and cannot be further split into separate propositions, and (3) whether the proposition is self-contained. The scores range from 1 to 3, where 1 means "no," 2 means "maybe," and 3 means "yes." We report the number of cases where the annotation was "no." The detailed instructions are provided in Table 8.
Appendix D Offline Indexing
Appendix E Retriever Models and QA Models
[…]
Appendix F Additional Results
In Section 5.2, we demonstrated the advantage of retrieval by proposition over retrieval by sentence, particularly as the population of the entity decreases in EQ. We used the occurrence in the top-1000 paragraphs retrieved by BM25 as a proxy for frequency, rather than counting the number of hyperlinks to the entity used in Sciavolino et al., 2021. Therefore, the trend in the performance versus frequency plot shows some differences (Figure 6) between our results and those in Sciavolino et al., 2021. For example, some entities are ambiguous (e.g., 1992, a TV series). In such cases, the occurrence of the surface form of the entity is large. Simultaneously, questions related to ambiguous entities are challenging to answer, leading to lower recall.
Appendix G Error Case Study
@@
To understand the source of errors from each type of retrieval granularity, we present and discuss four typical examples of mistakes in Table 6 and Table 7. With each example, we show the question and its corresponding top-1 retrieved text unit by the GTR retriever across the three granularities.
@@
We observe that with passage-level retrieval, the ambiguity of an entity or its references presents a challenge for dense retrievers, which echoes findings from Min et al. (2020). For instance, in example Q1, the question asks for “Super Bowl 50”, but the retrieved passage and sentence refers to “Super Bowl 5”. In Example Q2, passage retrieval fails to identify the part referring to the correct “atomic number”. Instead, the top-1 retrieved passage mentions “atomic number” in a different and irrelevant context to the question. Retrieval by sentences can also have a similar problem as retrieval by passages like Example Q1. Also, retrieval by sentences faces another challenge of lacking context. In Example Q3 (shown in Table 7), sentence-based retrieval fails as the correct sentence in the retrieved passage uses “it” to refer to the pericardial sac.
@@
Retrieval by propositions tackles the aforementioned problems by ensuring each retrieval unit contains one piece of fact only and necessary context is incorporated in the propositions. However, proposition-based retrieval faces challenges with questions that involve multi-hop reasoning over long-range textual analysis. In Example Q4 (shown in Table 7), the retrieved passage separately describes the actor’s name and the character they portray. There is not a single proposition that entails both the question and the answer.
@@
Passage Retrieval Sentence Retrieval Proposition Retrieval
@@
(a) Where was [X] born?
(b) Who was [X] created by?
@@
Figure 6: Document retrieval recall vs. the frequency of the target entity in each question from the Entity Questions dataset. We display the performance of two relations.
@@
Is the proposition fully supported by the passage? No : The information provided relates to the proposition, but there are some gaps or inconsistencies that prevent full support. Maybe : The information provided supports the proposition adequately, covering most aspects well; however, minor details or implications might not be fully explored or clarified. Yes: The information provided clearly and comprehensively addresses all aspects of the proposition, leaving no relevant details unexplained or ambiguous.
Should the given propositions be further split into separate propositions? No: The proposition has a compound structure that could be separated into distinct propositions. Maybe: The proposition is mostly straightforward with a single main idea and perhaps a minor additional detail. Splitting might enhance clarity but is not strictly necessary. Yes: The proposition is already concise and does not contain a compound structure. Splitting it into separate propositions would likely reduce clarity.
Is the given proposition self-contained? No: The proposition contains pronouns, terms, or references whose full names or meanings are not in the proposition. Maybe: The proposition is almost entirely self-contained, with only a few minor terms that might be ambiguous without additional context. Yes: The proposition is a self-contained claim without any ambiguities, fully understandable on its own.
Table 8: Instructions for data annotation in analyzing the quality of generated propositions.